Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Good Probability Question

Expert replies
by mjsobo » Tue Dec 08, 2009 9:36 am
Here is a fun and challenging probability question.

1) A box contains some Blue balls and 9 Red balls. If the probability that two balls randomly removed without replacement from the box are Red is 6/11, what is the total probability that the third ball randomly removed without replacement of the balls drawn is a blue ball?

The first step in answering this question is to find how many blue balls there are by solving for what we are given in the problem. We are given that the probability that two balls randomly removed without replacement from the box are red is 6/11, so then lets use this info to find x, the number of blue balls. We make an equation:

Let x be the number of blue balls
9/(9+x)*8/(8+x)=6/11 <--because there are first 9 balls, then 8 red balls.
Solving for x (skipping a few steps of basic algebra) we get x =3, as 9/12*8/11=6/11

Therefore we know that there are 3 balls. Now, because there are 3 blue balls left, and 7 red balls left, the odds of getting a blue ball next is 3/10. The next and final step is to multiply 6/11 by the probability of choosing a blue ball, or 6/11*3/10 = 9/55, as the question asks for total probability, given that the 1st 2 balls picked, without replacement, are red balls.
Good job!
Martin
GMAT Instructor with Grockit
Join the discussion
Source: — Problem Solving |